# Pretraining Data Detection for Large Language Models: A Divergence-based Calibration Method

### v2

@@ Abstract
As the scale of training corpora for large language models (LLMs) grows, model developers become increasingly reluctant to disclose details on their data. This lack of transparency poses challenges to scientific evaluation and ethical deployment. Recently, pretraining data detection approaches, which infer whether a given text was part of an LLM’s training data through black-box access, have been explored. The Min-KProb method, which has achieved state-of-the-art results, assumes that a non-training example tends to contain a few outlier words with low token probabilities. However, the effectiveness may be limited as it tends to misclassify non-training texts that contain many common words with high probabilities predicted by LLMs. To address this issue, we introduce a divergence-based calibration method, inspired by the divergence-from-randomness concept, to calibrate token probabilities for pretraining data detection. We compute the cross-entropy (i.e., the divergence) between the token probability distribution and the token frequency distribution to derive a detection score. We have developed a Chinese-language benchmark, PatentMIA, to assess the performance of detection approaches for LLMs on Chinese text. Experimental results on English-language benchmarks and PatentMIA demonstrate that our proposed method significantly outperforms existing methods.
@@ A Divergence-based Calibration Method
Weichao Zhang1,2,3 Ruqing Zhang1,2 Jiafeng Guo1,2thanks: Corresponding author Maarten de Rijke4 Yixing Fan1,2 Xueqi Cheng1,2 1CAS Key Lab of Network Data Science and Technology, ICT, CAS, Beijing, China. 2University of Chinese Academy of Sciences, Beijing, China. 3Zhongguancun Laboratory, Beijing, China. 4University of Amsterdam, Amsterdam, The Netherlands. {zhangweichao22z, zhangruqing, guojiafeng, fanyixing, cxq}@ict.ac.cn, m.derijke@uva.nl.
@@ 1 Introduction
The pretraining data detection problem can be viewed as an instance of the membership inference attack (MIA) task Shokri et al. (2017), where the primary objective is to determine if a particular text was part of a target LLM’s training corpus. Prevailing methods to tackle this problem are based on the idea that a text’s token probability distribution can reveal its inclusion in the training set. E.g., the Min-KProb method Shi et al. (2024) is based on the hypothesis that non-training examples tend to have more tokens assigned lower probabilities than training examples do. Min-KProb relies on the assumption that data with higher probability is more likely to be training data. Language models trained with a cross-entropy loss function tend to favor high-frequency tokens when conducting next-token prediction, which will also lead to LLMs generally predicting higher probabilities for high-frequency tokens Jiang et al. (2019). As illustrated in Figure 1(a), if the words of a non-training text are more common than those in a training text, it is more likely to be identified as training data under Min-KProb than the actual training text.
@@
We conduct experiments on two English-language benchmarks Shi et al. (2024) and on PatentMIA against a range of representative, state-of-the-art methods. Our experiments show that the proposed DC-rpPDD significantly outperforms prior methods. E.g., in the commonly used detection performance metrics, AUC and TPR@5%FPR, DC-PDD surpasses Min-KProb by $8.6$% and $13.3$%, respectively, on existing BookMIA benchmark.
@@ 5.2 Ablation studies
  • CLD: It serves as the initialization of DC-PDD by averaging all the calibrated token probabilities to compute $\beta$ for detection.
@@ 5.3 Impact of different factors
This section explores several factors that may influence the performance of DC-PDD, including two method-independent factors (model size and text length) and a method-dependent factor (the hyperparameter $a$).
@@ 7 Conclusion
DC-PDD, while showing promising results in pre-training data detection from LLMs, has several limitations. (i) DC-PDD utilizes a reference corpus to calculate the token frequency distribution to estimate that of the training corpus. Although working, the similarity between these two distributions remains uncertain. Additionally, the language of reference corpus should be the same as that of text to be detected. (ii) Secondly, an important hyperparameter in DC-PDD is the upper bound of calibrated token probabilities. We have demonstrated its significant impact on method performance, but not how the optimal value should be set. We leave this issue to future work. (iii) Thirdly, DC-PDD is specific to textual data. While some detection methods can be applied universally across different data modalities by relying on sample-level loss values obtained from models, our method is based on token-level probability. This specificity hinders its direct application to other types of data, such as images. (iv) Lastly, except for the closed-source model GPT-3 Brown et al. (2020), our research primarily focused on models with up to 20 billion parameters due to hardware constraints. Further studies replicating our work using larger-scale models will be essential to validate the effectiveness of DC-PDD in scenarios involving larger models.
Ethics Statement
@@ Acknowledgements
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62472408 and 62372431, the Strategic Priority Research Program of the CAS under Grants No. XDB0680102, XDB0680301, the National Key Research and Development Program of China under Grants No. 2023YFA1011602 and 2021QY1701, the Youth Innovation Promotion Association CAS under Grants No. 2021100, the Lenovo-CAS Joint Lab Youth Scientist Project, and the project under Grants No. JCKY2022130C039. This work was also (partially) funded by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl, project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21, which is (partly) financed by the Dutch Research Council (NWO), and the FINDHR (Fairness and Intersectional NonDiscrimination in Human Recommendation) project that received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.
@@ A.1 Baseline details
Zlib. Carlini et al. (2021) Similarly to the Small Ref method, but use the zlib entropy of $x$ in place of the smaller model’s perplexity for $x$. The zlib entropy is the entropy in bits when the sequence is compressed using zlib  [^3] https://github.com/madler/zlib. The detection score is then determined by the ratio of $\mathcal{M}$’s perplexity for $x$ to the zlib entropy of $x$. A low score suggests that $x$ was likely part of the pretraining data.

### v3

@@ Abstract
As the scale of training corpora for large language models (LLMs) grows, model developers become increasingly reluctant to disclose details on their data. This lack of transparency poses challenges to scientific evaluation and ethical deployment. Recently, pretraining data detection approaches, which infer whether a given text was part of an LLM’s training data through black-box access, have been explored. The Min-KProb method, which has achieved state-of-the-art results, assumes that a non-training example tends to contain a few outlier words with low token probabilities. However, the effectiveness may be limited as it tends to misclassify non-training texts that contain many common words with high probabilities predicted by LLMs. To address this issue, we introduce a divergence-based calibration method, inspired by the divergence-from-randomness concept, to calibrate token probabilities for pretraining data detection. We compute the cross-entropy (i.e., the divergence) between the token probability distribution and the token frequency distribution to derive a detection score. We have developed a Chinese-language benchmark, PatentMIA, to assess the performance of detection approaches for LLMs on Chinese text. Experimental results on English-language benchmarks and PatentMIA demonstrate that our proposed method significantly outperforms existing methods. Our code and PatentMIA benchmark are available at https://github.com/zhang-wei-chao/DC-PDD.
@@ A Divergence-based Calibration Method
Weichao Zhang1,2,3 Ruqing Zhang1,2 Jiafeng Guo1,2thanks: Corresponding author Maarten de Rijke4 Yixing Fan1,2 Xueqi Cheng1,2 1CAS Key Lab of Network Data Science and Technology, ICT, CAS, Beijing, China 2University of Chinese Academy of Sciences, Beijing, China 3Zhongguancun Laboratory, Beijing, China 4University of Amsterdam, Amsterdam, The Netherlands {zhangweichao22z, zhangruqing, guojiafeng, fanyixing, cxq}@ict.ac.cn, m.derijke@uva.nl
@@ 1 Introduction
The pretraining data detection problem can be viewed as an instance of the membership inference attack (MIA) task Shokri et al. (2017), where the primary objective is to determine if a particular text was part of a target LLM’s training corpus. Prevailing methods to tackle this problem are based on the idea that a text’s token probability distribution can reveal its inclusion in the training set. E.g., the Min-KProb method Shi et al. (2024) is based on the hypothesis that non-training examples tend to have more tokens assigned lower probabilities than training examples do. Min-KProb relies on the assumption that data with higher probability is more likely to be training data. Language models trained with a cross-entropy loss function tend to favor high-frequency tokens when conducting next-token prediction, which will also lead to LLMs generally predicting higher probabilities for high-frequency tokens Jiang et al. (2019). As the conceptual example shown in figure 1, $x^{1}$ is a non-training text and $x^{2}$ is a training text. We can see that the lowest raw token probabilities for $x^{1}$ are higher than those for $x^{2}$, which may be because the words in $x^{1}$ (e.g., "boys", "great") are generally more common than the words in $x^{2}$ (e.g., "erudite", "conundrum"). Therefore, Min-kProb will calculate a detection score of -0.88 for $x^{1}$ and -2.94 for $x^{1}$, which means that $x^{1}$ is more likely to be considered a training text than. This is contrary to the actual situation.
@@
We conduct experiments on two English-language benchmarks Shi et al. (2024) and on PatentMIA against a range of representative, state-of-the-art methods. Our experiments show that the proposed DC-PDD significantly outperforms prior methods. E.g., in the commonly used detection performance metrics, AUC and TPR@5%FPR, DC-PDD surpasses Min-KProb by $8.6$% and $13.3$%, respectively, on existing BookMIA benchmark.
@@ 5.2 Ablation Studies
  • CLD: It serves as the initialization of DC-PDD by averaging all the CaLibrateD token probabilities to compute a score for detection.
@@ 5.3 Impact of Different Factors
This section explores several factors that may influence the performance of DC-PDD, including two method-independent factors (model size and text length) and two method-dependent factor (the reference corpus $\mathcal{D^{\prime}}$ and hyperparameter $a$).
@@
Reference corpus $\mathcal{D^{\prime}}$. Recall that we use a reference corpus $\mathcal{D^{\prime}}$ to estimate the token frequency distribution of the LLM’s pretraining corpus, w.r.t. Eq. (4). To analyze the effect of different reference corpora on the efficacy of the method, we compare the performance of DC-PDD under various reference corpus settings across different scales and domains. Specifically, when detecting WikiMIA-128 from pythia-6.9B, we employee $\approx 1$Gb of C4 corpus, $\approx 10$Gb of C4 corpus, $\approx 1$Gb of Case-law corpus, and $\approx 10$Gb of Case-law corpus as the reference corpus respectively. Note that the Case-law Louis Brulé Naudet (2024) is a corpus in the legal domain. As shown in Table 4, We observe that the performance of DC-PDD does not exhibit significant differences across the various reference corpora, indicating that DC-PDD is not sensitive to the selection of a reference corpus. Notably, when the reference corpus is chosen as the $\approx 10$Gb of C4 corpus, the performance of DC-PDD is the best. This enhancement may be attributed to the greater diversity of the C4 corpus compared to the $\approx 10$Gb of Case-law corpus, as well as the richer data compared to the $\approx 1$Gb of C4 corpus, which allow for a more accurate estimation of the token frequency distribution in the LLM’s pretraining corpus, thereby resulting in better performance.
$\mathcal{D^{\prime}}$ C4 Case-law
$\approx 1$Gb $\approx 10$Gb $\approx 1$Gb $\approx 10$Gb
Pythia-6.9B 0.688 0.698 0.687 0.688
Table 4: AUC scores of DC-PDD in different reference corpus settings.
@@ 7 Conclusion
DC-PDD, while showing promising results in pre-training data detection from LLMs, has several limitations. (i) DC-PDD utilizes a reference corpus to calculate the token frequency distribution to estimate that of the training corpus. Although working, the similarity between these two distributions remains uncertain. Additionally, the language of reference corpus should be the same as that of text to be detected. (ii) Secondly, an important hyperparameter in DC-PDD is the upper bound of calibrated token probabilities. We have demonstrated its significant impact on method performance, but not how the optimal value should be set. We leave this issue to future work. (iii) Thirdly, DC-PDD is specific to textual data. While some detection methods can be applied universally across different data modalities by relying on sample-level loss values obtained from models, our method is based on token-level probability. This specificity hinders its direct application to other types of data, such as images. (iv) Fourthly, DC-PDD requires access to token probabilities, and therefore is not applicable to some closed-source models. In the future, we will explore detection methods based solely on model output to design more generalizable detection methods. (v) Lastly, except for the closed-source model GPT-3 Brown et al. (2020), our research primarily focused on models with up to 20 billion parameters due to hardware constraints. Further studies replicating our work using larger-scale models will be essential to validate the effectiveness of DC-PDD in scenarios involving larger models.
Ethical Considerations
@@ Acknowledgements
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62472408 and 62372431, the Strategic Priority Research Program of the CAS under Grants No. XDB0680102 and XDB0680301, the National Key Research and Development Program of China under Grants No. 2023YFA1011602 and 2021QY1701, the Youth Innovation Promotion Association CAS under Grants No. 2021100, the Lenovo-CAS Joint Lab Youth Scientist Project, and the project under Grants No. JCKY2022130C039. This work was also (partially) funded by the Dutch Research Council (NWO), under project numbers 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006, and the European Union’s Horizon Europe program under grant agreement No 101070212.All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.
@@ A.1 Baseline details
Zlib. Carlini et al. (2021) Similar to the Small Ref method, but uses the zlib entropy of $x$ in place of the smaller model’s perplexity for $x$. The zlib entropy is the entropy in bits when the sequence is compressed using zlib. [^3] https://github.com/madler/zlib The detection score is then determined by the ratio of $\mathcal{M}$’s perplexity for $x$ to the zlib entropy of $x$. A low score suggests that $x$ was likely part of the pretraining data.